OpenStudy (anonymous):

Help!!! what is the magnitude of the acceleration of a speck of clay on the edge of a potters wheel turning at 45.0rpm (revolutions per minute) if its diameter is 30.0cm?

OpenStudy (shamim):

Given a=? n=45/60 hz r=45/2cm=22.5cm=.225m

OpenStudy (shamim):

U hv to use the formula a=omega^2*r Omega=2*pi*n

OpenStudy (shamim):

Feel free to ask for more clarification